OpenStudy (anonymous):

Solve for x 3 = 2 +5e^-4x

OpenStudy (anonymous):

\[3 = 2+5e ^{-4x}\]

OpenStudy (anonymous):

How about you move that 2 over to the other side

OpenStudy (anonymous):

\[1 =5e ^{-4x}\]

OpenStudy (anonymous):

Then, divide over that 5: \[1/5 = 5e ^{-4x}/5\]

OpenStudy (anonymous):

So then we end up getting: \[\frac{1}{5} = e^{-4x}\]

OpenStudy (anonymous):

Now, we want to get rid of that "e". How can we do so?

OpenStudy (anonymous):

What do we know "cancels" out the base "e"?

OpenStudy (anonymous):

Is it using ln?

OpenStudy (anonymous):

Yes! Good job

OpenStudy (anonymous):

So now, we ln both sides: \[\ln(\frac{1}{5}) = \ln(e^{-4x})\]

OpenStudy (anonymous):

So what happens to the ln(e..) expression?

OpenStudy (anonymous):

They essentially "cancel" each other out, leaving you with ____

OpenStudy (anonymous):

Understand? Do you know how to carry on from here?

OpenStudy (anonymous):

I divide it by 4?

OpenStudy (anonymous):

Yes, you should end up getting (for the right hand side), a -4x remaining after you "cancel" out the ln and "e".

OpenStudy (anonymous):

No, you divide by NEGATIVE 4

OpenStudy (anonymous):

And then you should be done, unless you want an estimate. The value: \[\frac{\ln(\frac{1}{5})}{-4}\] is an exact value.

OpenStudy (anonymous):

So if you want the exact answer, that would be it.

OpenStudy (anonymous):

Wow, thank you. I keep forgetting how and when to use the formulas. Thanks again!

OpenStudy (anonymous):

So the idea of these types of problems is that you: 1) want to get rid of the e's

OpenStudy (anonymous):

This is what you should do when experiencing an equation to solve for x: ask yourself, what is the problem, and how can I get rid of it?

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

Your one of the best people explaining equations on here! Thanks for helping me out and giving me a tip! :)

OpenStudy (anonymous):

I am sure there are others too that are far more knowledgable in this board than me, wait till they help you out. :)